Hermin is 6 years older than Doy. Eight years from now, Hermin will be 4 times as old as Doy was 4 years ago. Find their present ages?
Answer by ptaylor(2198)

Let x=Doy's present age; x+8=Doy's age 8 years from now; x-4=Doy's age 4 years ago
Then x+6=Hermin's present age; x+14=Hermin's age 8 years from now
Now we are told the following:
x+14=4(x-4) get rid of parens (distributive)
x+14=4x-16 subtract x from and add 16 to each side
x-x+14+16=4x-x-16+16 collect like terms
30=3x divide each side by 3
x=10-----------------------Doy's present age
x+6=10+6=16 ----------------Hermin's present age
CK
16+8=4(10-4)
24=24
Does this help???----------------ptaylor